### v1

@@ 1 Introduction
While BNS mergers are the only class of GW event with a confirmed EM counterpart, it is also predicted that binary black hole (BBH) mergers can produce counterparts in certain circumstances. The majority of proposed counterpart mechanisms involve the interaction of the binary or merger remnant with a coexistent medium, usually the gaseous accretion disk of an active galactic nucleus (AGN), whether through accretion (Bartos et al., 2017b), ram-pressure stripping of gas about the remnant or jetted Bondi accretion (McKernan et al., 2019), or breakout emission from a post-merger jet (Tagawa et al., 2023, 2024). A cartoon of the AGN-associated EM counterpart mechanism is show in Figure 1.
@@ 2.1 GW data
Higher mass () BBH mergers are more likely to have originated from the disks of AGNs (e.g. Gayathri et al. 2021), and so we also inform our decision to trigger BBH follow up by estimating the total mass of the binary. We follow a similar calculation to that in Graham et al. (2020), and consider that (Berry et al., 2015), as well as (Finn & Chernoff, 1993), where is the chirp mass of the binary. For this event, we assume a 1.4-1.4 binary neutron star merger detection horizon of 150 and 152 Mpc for Hanford and Livingston, respectively, an equal mass system, and the luminosity distance marginalized over the sky; we derive a total rest frame mass of for S230922g. Note that this estimate is highly uncertain with an error bar of at least a factor of 2.
@@ 2.2 DECam observations
Leading BBH counterpart models predict delays of 1-3 weeks after the merger before the manifestation of any EM counterpart, due to the need for the signal to first emerge from the depths of the AGN disk. Even so, it is valuable to initiate EM follow up for these kinds of events within a few days of the GW trigger, as high-cadence archival data are generally not available in the localization area of a given GW event and a baseline of sources in the localization area can be useful when vetting candidates. Accordingly, we decided to trigger our Target of Opportunity (ToO) program for the night following the event (the evening of September 22, 2023) in order to establish a baseline of activity at the GW merger time and to facilitate the identification of novel phenomena in subsequent observing epochs.
@@ 3.1 Automated vetting
1 real/bogus score 0.9 | 17,515 | 0.075 |
2 detections with min | 3558 | 0.015 |
@@
We first remove any detections on difference images contaminated by bad pixels recorded in the DECam CP data quality mask products. Surviving features were then scored using a Convolutional Neural Network (CNN) trained on archival DECam data products from our pipeline to perform real/bogus classification. Possible CNN scores range from 0 to 1, with higher scores associated with more realistic astrophysical objects; at this step we kept any detections with scores greater than 0.7. Roughly 6% of the initial list of detections passed these steps to form our list of real astrophysical detections.
@@ 3.2 Additional vetting
AT 2023uab | 0.128 | Specz (this work) | 0.785 | 0.554 | SNIa | 0.814 |
@@ 4.1 AT 2023aagj (C2023092422006400m275139)
Wide-field Infrared Survey Explorer (WISE) (Wright et al., 2010a) infrared data for the host of this event is shown in Figure 3. Following Clark et al. (2024), we retrieved this data from the NASA/IPAC infrared science archive (IRSA)111111https://irsa.ipac.caltech.edu/, utilizing data from both the ALLWISE (Wright et al., 2010b) and NEOWISE Reactivation Releases (NEOWISE-R; Mainzer et al., 2011, 2014). Individual observations are combined following quality control cuts to obtain one mean magnitude per filter for each observation period, providing a 6 month cadence. Notably, there are variations in the host system of similar amplitude to our detected transient, but it must be recognized that such variations occur on much longer () timescales than the timescale of AT 2023aagj. We compare this behaviour with publicly available PanSTARRS (Flewelling et al., 2020) and DECam images at this location, which, although sparse, do not show any significant change in flux. We observe that a significant jump of mag in band 5” aperture magnitude occurred according to the SkyMapper data release (Onken et al., 2024) between MJD 57650 and 57850. It is possible that an associated optical signature existed in that time period, but was missed by the relatively sparse nature of the optical observations, and the infrared (IR) component we observed was delayed from the optical via dust reprocessing. We do not observe a rebrightening of the source in the WISE bands over the timescale of the optical transient, but there appears to be a increase in magnitude in December 2023, potentially indicating the start of a reprocessed, delayed emission from the event.
@@ 5 Discussion
Taking our favored counterpart candidate of AT 2023aagj, we work under the assumption that the transient is related to S230922g and try to infer whether the the properties of the flare may be consistent with expectations for the counterpart signal, mechanism, and host.
What can be noted is that the persistent blue color is unlike typical supernovae, and structure function (SF) arguments as in Palmese et al. (2021) show that this transient is distinguished from typical AGN activity at the 5.9 level (corresponding to a one-sided probability ). Taking the 90% CI volume of the event as Mpc3 (as calculated with ligo.skymap) and a fiducial AGN number density of Mpc-3 (Bartos et al., 2017a; Greene & Ho, 2007), we estimate a count of 6600 AGN in the volume. Alternatively, crossmatching the Milliquas (Flesch, 2023) quasar catalog with the skymap (also with ligo.skymap) yields a 3D match of 826 objects. Given the discovery window of our follow-up campaign days (calculated as the difference in time between the second epoch and the last), we calculate the probability of a chance discovery of a similar flare as
which predicts and for the number density and Milliquas estimates, respectively.
@@
If we assume from Figure 3 that the flare begins 2 days post-S230922g (denoted by a vertical line), and the flare lasts through the spectra denoted by the blue and red lines, then the overall flare duration is days. Given , the average luminosity of the flare is , which can be parameterized as where is the Eddington luminosity. If we parameterize the Bondi accretion rate onto the merged BH () as:
@@ Acknowledgements
TC, AP, and LH acknowledge that this material is based upon work supported by NSF Grant No. 2308193. BO gratefully acknowledges support from the McWilliams Postdoctoral Fellowship at Carnegie Mellon University. BM & KESF are supported by NSF AST-2206096 and NSF AST-1831415 and Simons Foundation Grant 533845 as well as Simons Foundation sabbatical support. The Flatiron Institute is supported by the Simons Foundation. AP thanks Rosalba Perna, Armin Rest, and Stephen Smartt for useful discussion.
@@ Appendix A Additional data
For some transients we took additional data with the Three Channel Camera (3KK) (Lang-Bardl et al., 2016) at the 2.1m-Fraunhofer Wendelstein Telescope (Hopp et al., 2014) located in the German alps. The instrument is capable of taking three images at the same time in two optical passbands and one near-infrareds (NIRs) band. For the monitoring of the transients we used the , and bands. The data were reduced with a pipeline developed at the University Observatory Munich. For the detrending of the images the pipeline makes use of the tools from Gössl & Riffeser (2002). Then the images are calibrated and coadded using SExtractor (Bertin & Arnouts, 1996), SCAMP (Bertin, 2006) and SWarp (Bertin et al., 2002), difference photometry is conducted with SFFT (Hu et al., 2022), and extinction corrections are applied as described in §3.1. During this process, the astrometric solution is computed against the Gaia EDR3 catalog (Gaia Collaboration et al., 2016, 2021) and the zero points for the optical bands are calibrated with the PanSTARRS 1s (PS1s) catalog (Magnier et al., 2013). The NIRs zero points are matched to the Two Micron All Sky Surveys (2MASSs) catalog (Skrutskie et al., 2006) and converted to AB magnitudes (Blanton & Roweis, 2007). Wendelstein data are plotted as outlined diamonds in Figure 5, with , , and band data plotted in green, yellow, and red, respectively.

### v2

@@ 1 Introduction
While BNS mergers are the only class of GW event with a confirmed EM counterpart, it is also predicted that binary black hole (BBH) mergers can produce counterparts in certain circumstances. The majority of proposed counterpart mechanisms involve the interaction of the binary or merger remnant with a coexistent medium, usually the gaseous accretion disk of an active galactic nucleus (AGN), whether through accretion (Bartos et al., 2017b), ram-pressure stripping of gas about the remnant or jetted Bondi accretion (McKernan et al., 2019), or breakout emission from accretion (Kimura et al., 2021; Rodríguez-Ramírez et al., 2024) or a post-merger jet (Tagawa et al., 2023, 2024). A cartoon of the AGN-associated EM counterpart mechanism is show in Figure 1.
@@ 2.1 GW data
Higher mass () BBH mergers are more likely to have originated from the disks of AGNs (e.g. Gayathri et al. 2021), and so we also inform our decision to trigger BBH follow up by estimating the total mass of the binary. We follow a similar calculation to that in Graham et al. (2020), and consider that (Berry et al., 2015), as well as (Finn & Chernoff, 1993), where is the chirp mass of the binary. For this event, we assume a 1.4-1.4 binary neutron star merger detection horizon of 150 and 152 Mpc for Hanford and Livingston, respectively, an equal mass system, and the luminosity distance marginalized over the sky; we derive a total rest frame mass of for S230922g. The BNS merger detection horizon serves as an additional scaling factor in the overall calculation. Note that this estimate is highly uncertain with an error bar of at least a factor of 2. This uncertainty is informed from the calibration of the proposed mass relationship with the LIGO O3 BBH data set.
@@ 2.2 DECam observations
Because most BBH counterpart models concern the emergence of an EM signature from the interior of an AGN disc, delays of days are expected before such a signature becomes observable. Even so, it is valuable to initiate EM follow up for these kinds of events within a few days of the GW trigger, as high-cadence archival data are generally not available in the localization area of a given GW event and a baseline of sources in the localization area can be useful when vetting candidates. Accordingly, we decided to trigger our Target of Opportunity (ToO) program for the night following the event (the evening of September 22, 2023) in order to establish a baseline of activity at the GW merger time and to facilitate the identification of novel phenomena in subsequent observing epochs.
@@ 3.1 Automated vetting
| 1 real/bogus score 0.9 | 17,515 | 0.075 |
| 2 detections with min | 3558 | 0.015 |
@@
We first remove any detections on difference images contaminated by bad pixels recorded in the DECam CP data quality mask products. Surviving features were then scored using a rotation-invariant convolutional neural network (CNN) (Cabrera-Vives et al., 2017; Sun et al., 2022) trained on archival DECam data products from our pipeline to perform real/bogus classification. Possible CNN scores range from 0 to 1, with higher scores associated with more realistic astrophysical objects. From testing on archival DECam data a threshold of 0.7 was found to facilitate a 98% recall rate with a 7.2% bogus contamination rate; this threshold was accepted as our cutoff for this analysis. Roughly 6% of the initial list of detections passed these steps to form our list of real astrophysical detections.
@@ 3.2 Additional vetting
| AT 2023uab | 0.128 | Specz (this work) | 0.785 | 0.554 | SNIa | 0.814 | |
@@ 4.1 AT 2023aagj (C202309242206400m275139)
Wide-field Infrared Survey Explorer (WISE) (Wright et al., 2010a) infrared data for the host of this event is shown in Figure 3. Following Clark et al. (2024), we retrieved this data from the NASA/IPAC infrared science archive (IRSA)111111https://irsa.ipac.caltech.edu/, utilizing data from both the ALLWISE (Wright et al., 2010b) and NEOWISE Reactivation Releases (NEOWISE-R; Mainzer et al., 2011, 2014). Individual observations are combined following quality control cuts to obtain one mean magnitude per filter for each observation period, providing a 6 month cadence. The host is found to have brightened by 0.2 mag (with typical photometric uncertainties 0.05 mag) in both W1 and W2 bands starting around MJD 57900. The SkyMapper data release (Onken et al., 2024) shows that the object brightened by 0.13/0.1 mag (uncertainties 0.05 mag) in /-band 5” aperture photometry over MJD 57870-57875. A possible explanation for these phenomena identifies the infrared (IR) component as a reprocessed signature of the optical event; while blue, this event is slower than AT 2023aagj, which at one point brightened by 0.4 mag in both and band over a 5 day interval. We also examine archival DECam and PanSTARRS (Flewelling et al., 2020) data covering this object, but are unable to put further constraints on past activity in part due to the sparse temporal coverage of these datasets in this region.
@@ 5 Discussion
We return to our favored counterpart candidate AT 2023aagj and assess its significance as an AGN transient. The persistent blue color of the transient distinguishes it from the phenomenology of typical supernovae, a separation further supported by a ParSNIP classification of “TDE”/“Non-SN” as noted above. We apply structure function (SF) arguments as presented in (Palmese et al., 2021) to the photometry to calculate the probability that the observed variation occurred as a result of typical AGN activity. We find the variability of AT 2023aagj significant on the 5.9 level, with an associated one-sided Gaussian probability of ; such a low value makes it highly unlikely that AT 2023aagj is a feature of the host AGN’s activity, and strongly prefers the model of AT 2023aagj as a specially transient phenomenon. Taking the 90% CI volume of the event as Mpc3 (as calculated with ligo.skymap) and fiducial AGN number densities of Mpc-3 and Mpc-3 (Bartos et al. 2017a; Greene & Ho 2007; the two values reflect expected and conservative (more AGN/background flares) rates), we estimate counts of 6600 and 37000 AGN in the volume. Alternatively, crossmatching the Milliquas (Flesch, 2023) quasar catalog with the skymap (also with ligo.skymap) yields a 3D match of 826 objects. Given the discovery window of our follow-up campaign days (calculated as the difference in time between the second epoch and the last), we calculate the probability of a chance discovery of a similar flare as
which predicts and , and for the number density and Milliquas estimates, respectively. We note that while these rates distinguish AT 2023aagj as a special event, due to uncertainties in the understanding of BBH counterpart morphology the identity of AT 2023aagj as a counterpart to S230922g remains largely uncertain. A caveat of this analysis is that we also cannot account for the variability of this specific AGN given the lack of long-term monitoring from archival data.
@@
If we assume from Figure 3 that the flare begins 2 days post-S230922g, and the flare lasts through the spectra denoted by the blue and red lines, then the overall flare duration is days. We note that this is to be considered a rough estimate of the flare timescale in what follows. Given , the average luminosity of the flare is , which can be parameterized as where is the Eddington luminosity. If we parameterize the Bondi accretion rate onto the merged BH () as:
@@ 6 Conclusions
Data and Code Availability
The data products and figure code for this publication are available at https://doi.org/10.5281/zenodo.13787730.
@@ Acknowledgements
TC, AP, and LH acknowledge that this material is based upon work supported by NSF Grant No. 2308193. BO gratefully acknowledges support from the McWilliams Postdoctoral Fellowship at Carnegie Mellon University. BM & KESF are supported by NSF AST-2206096 and NSF AST-1831415 and Simons Foundation Grant 533845 as well as Simons Foundation sabbatical support. The Flatiron Institute is supported by the Simons Foundation. AP thanks Rosalba Perna, Armin Rest, and Stephen Smartt for useful discussion. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311.
@@ Appendix A Additional data
For some transients we collected additional data with the Three Channel Camera (3KK) (Lang-Bardl et al., 2016) at the 2.1m-Fraunhofer Wendelstein Telescope (Hopp et al., 2014) located in the German alps. The instrument takes three images at the same time in two optical passbands and one near-infrared (NIR) band. For the monitoring of the transients we used the , and bands. For the detrending of the images the pipeline makes use of the tools from Gössl & Riffeser (2002). Then the images are calibrated and coadded using SExtractor (Bertin & Arnouts, 1996), SCAMP (Bertin, 2006) and SWarp (Bertin et al., 2002), difference photometry is conducted with SFFT (Hu et al., 2022), and extinction corrections are applied as described in §3.1. During this process, the astrometric solution is computed against the Gaia EDR3 catalog (Gaia Collaboration et al., 2016, 2021) and the zero points for the optical bands are calibrated with the PanSTARRS 1s (PS1s) catalog (Magnier et al., 2013). The NIRs zero points are matched to the Two Micron All Sky Surveys (2MASSs) catalog (Skrutskie et al., 2006) and converted to AB magnitudes (Blanton & Roweis, 2007). Wendelstein data are plotted as outlined diamonds in Figure 5, with , , and band data plotted in green, yellow, and red, respectively.